# Bayesian Density-Density Regression with Application to Cell-Cell Communications

### v1

@@ 3.1 Generalized Bayesian Density-Density Regression
where $f$ is a measurable function that maps from $\mathbb{R}^{d_{1}}$ to $\mathbb{R}^{d_{2}}$, $L\geq 1$, and $w>0$. The likelihood $p(G_{i}\mid F_{i},f)$ in (5) is a generalized likelihood (Bissiri et al., , 2016).
@@
In (5) we are using SW distance to define a loss function that serves as generalized (negative log) likelihood. SW is an attractive choice for the loss function as it naturally defines a distance between random distributions $f\sharp{F}_{i}$ and ${G}_{i}$. One of the advantages of using the generalized likelihood is that it sidesteps the need for a full probabilistic description of all relevant unknowns. If desired, the SW distance between two empirical distributions can be replaced by the SW distance between any alternative estimates of $F_{i}$ and $G_{i}$, for example, using kernel density estimates or a fitted mixture of Gaussians. For simplicity, we choose to use empirical distributions. Also, SW distance, which relies on Euclidean geometry, can be replaced with other variants of SW (Kolouri et al., , 2019) if more information about the geometry of the distributions were available.
@@
To assess goodness of fit, we use the following predictive error between the fitted distribution and the recorded response distributions: $PE(\mathcal{S},L)=\frac{1}{N}\sum_{i=1}^{N}\widehat{SW}_{2}^{2}(f\sharp{F}_{i},{G}_{i};L),$ which can be calculated by posterior samples. Given $f_{1},\ldots,f_{T}\sim p(f\mid\mathcal{S})$, we have the following approximated mean predictive error: $\overline{\text{PE}}(\mathcal{S},L)=\frac{1}{T}\frac{1}{N}\sum_{t=1}^{T}\sum_{i=1}^{N}\widehat{SW}_{2}^{2}(f_{t}\sharp{F}_{i},{G}_{i};L).$ Assume we have a reference model $f_{0}$ that serves as a worst-case reference (see later for examples), which is similar to an intercept-only model in a linear regression. We normalize the relative error to obtain the following relative predictive error (RPE):
$\text{RPE}(\mathcal{S},L)=\frac{1}{N}\sum_{i=1}^{N}\frac{\widehat{SW}_{2}^{2}(f\sharp{F}_{i},{G}_{i};L)}{\widehat{SW}_{2}^{2}(f_{0}\sharp{F}_{i},{G}_{i};L)}.$ (6)
@@
Similarly, the mean RPE can be estimated using posterior samples, i.e.,
$\overline{\text{RPE}}(\mathcal{S},L)=\frac{1}{T}\frac{1}{N}\sum_{t=1}^{T}\sum_{i=1}^{N}\frac{\widehat{SW}_{2}^{2}(f_{t}\sharp{F}_{i},{G}_{i};L)}{\widehat{SW}_{2}^{2}(f_{0t}\sharp{F}_{i},{G}_{i};L)},$ (7)
@@
3.2 Posterior Inference
Posterior inference under (5) is based on the generalized posterior,
$p(f\mid\mathcal{S})\propto p(f)\exp\left(-w\sum_{i=1}^{N}\widehat{SW}_{2}^{2}(f\sharp{F}_{i},{G}_{i};L)\right),$ (8)
where $p(f)$ is a prior distribution of $f$. We implement inference by Markov chain Monte Carlo (MCMC) simulation from (8). For computational efficiency, we restrict the function $f$ to a parametric form, i.e., $f_{\phi}$ with parameters $\phi\in\Phi$. Given a prior $p(\phi)$, we have the generalized posterior:
$p(\phi\mid\mathcal{S})\propto p(\phi)\exp\left(-w\sum_{i=1}^{N}\widehat{SW}_{2}^{2}(f_{\phi}\sharp{F}_{i},{G}_{i};L)\right).$ (9)
@@ 3.2 Posterior Inference
We restrict the function $f_{\phi}$ and $\log\{p(\phi)\}$ to be differentiable with respect to $\phi$ to facilitate the use of gradient-based MALA (one-step hybrid Monte Carlo) transition probabilities using the following proposal distribution: $q(\phi^{\prime}\mid\phi)\propto\exp\left(-\frac{1}{4\eta}\|\phi^{\prime}-\phi-\eta\nabla_{\phi}\log p(\phi\mid\mathcal{S})\|_{2}^{2}\right)$, for a fixed step size $\eta>0$. We sample $\phi^{\prime}\sim q(\phi^{\prime}\mid\phi)$ as $\phi^{\prime}=\phi+\eta\nabla_{\phi}\log p(\phi\mid\mathcal{S})+\sqrt{2\eta}\epsilon_{0}$, with $\epsilon_{0}\sim\mathcal{N}(0,I)$, a standard multivariate Gaussian distribution of dimension matching the parameter $\phi$. We accept $\phi^{\prime}$ with probability: $\alpha=1\wedge{p(\phi\mid\mathcal{S})q(\phi\mid\phi^{\prime})}\big{/}{\left(p(\phi^{\prime}\mid\mathcal{S})q(\phi^{\prime}\mid\phi)\right)}.$
@@
We now derive the gradient $\nabla_{\phi}\log p(\phi\mid\mathcal{S})=\nabla_{\phi}\log p(\phi)-w\sum_{i=1}^{N}\nabla_{\phi}SW_{p}^{p}(f_{\phi}\sharp{F}_{i},G_{i})$, where $\nabla_{\phi}\widehat{SW}_{2}^{2}(f_{\phi}\sharp{F}_{i},{G}_{i};L)=\frac{1}{L}\sum_{l=1}^{L}\nabla_{\phi}W_{2}^{2}(\theta_{l}\sharp f_{\phi}\sharp{F}_{i},\theta_{l}\sharp{G}_{i})$ with
@@
$\nabla_{\phi}W_{2}^{2}(\theta_{l}\sharp f_{\phi}\sharp{F}_{i},\theta_{l}\sharp{G}_{i})=\nabla_{\phi}\left(\min_{\gamma\in\Gamma(\mathbf{1}/m_{i1},\mathbf{1}/m_{i2})}\sum_{i^{\prime}=1}^{m_{i1}}\sum_{j^{\prime}=1}^{m_{i2}}(\theta_{l}^{\top}f_{\phi}(x_{ii^{\prime}})-\theta_{l}^{\top}y_{ij^{\prime}})^{2}\gamma_{i^{\prime}j^{\prime}}\right).$ (10)
@@
$\nabla_{\phi}\left(\min_{\gamma\in\Gamma\left(\frac{\mathbf{1}}{m_{i1}},\frac{\mathbf{1}}{m_{i2}}\right)}\sum_{i^{\prime}=1}^{m_{i1}}\sum_{j^{\prime}=1}^{m_{i2}}(\theta^{\top}f_{\phi}(x_{ii^{\prime}})-\theta^{\top}y_{ij^{\prime}})^{2}\gamma_{i^{\prime}j^{\prime}}\right)=\nabla_{\phi}\left(\sum_{i^{\prime}=1}^{m_{i1}}\sum_{j^{\prime}=1}^{m_{i2}}(\theta^{\top}f_{\phi}(x_{ii^{\prime}})-\theta^{\top}y_{ij^{\prime}})^{2}\gamma_{i^{\prime}j^{\prime}}^{\star}\right)\\ =\sum_{i^{\prime}=1}^{m_{i1}}\sum_{j^{\prime}=1}^{m_{i2}}\gamma_{i^{\prime}j^{\prime}}^{\star}\nabla_{\phi}(\theta^{\top}f_{\phi}(x_{ii^{\prime}})-\theta^{\top}y_{ij^{\prime}})^{2}\\ =2\sum_{i^{\prime}=1}^{m_{i1}}\sum_{j^{\prime}=1}^{m_{i2}}\gamma_{i^{\prime}j^{\prime}}^{\star}(\theta^{\top}f_{\phi}(x_{ii^{\prime}})-\theta^{\top}y_{ij^{\prime}})\nabla_{\phi}(\theta^{\top}f_{\phi}(x_{ii^{\prime}})).$ (11)
The calculation of $\nabla_{\phi}(\theta^{\top}f_{\phi}(x_{ii^{\prime}}))$ depends on the form of $f_{\phi}$. In the upcoming examples and simulations, we utilize a linear function for simplicity, $f_{A,b}(x)=Ax+b$, i.e., $\phi=(A,b)$ with $A=[A_{ij}]\in\mathbb{R}^{d_{2}\times d_{1}}$ and $b\in\mathbb{R}^{d_{2}}$. We put a horseshoe prior (Carvalho et al., , 2009) on $A$, which will be later discussed in detail in the context of graph discovery:
@@
and $b\sim\mathcal{N}(0,I)$. We implement posterior simulation using standard MCMC transition probabilities using the inverse gamma complete conditional posterior distributions for $\lambda_{ij}^{2}$, $\nu_{ij}$, $\tau^{2}$, and $\zeta$. We refer the reader to Supplementary Materials  A for the detail of the MCMC sampler. To sample from $p(A,b\mid\lambda^{2},\tau^{2},\mathcal{S})$, we use the MALA algorithm as discussed above. Plugging in $\nabla_{A}(\theta^{\top}(Ax_{ii^{\prime}}+b))=\theta x_{ii^{\prime}}^{\top}$ and $\nabla_{b}(\theta^{\top}(Ax_{ii^{\prime}}+b))=\theta$ into Equation (11), we obtain the gradient for the log-likelihood. For the gradient of the log-prior, we have $\nabla_{A_{ij}}\log p(A_{ij}\mid\lambda_{ij}^{2},\tau^{2})=\nabla_{A_{ij}}\frac{-A_{ij}^{2}}{2\lambda_{ij}^{2}\tau^{2}}=\frac{-A_{ij}}{\lambda_{ij}^{2}\tau^{2}}$ and $\nabla_{b}\log p(b)=\nabla_{b}\frac{-\|b\|_{2}^{2}}{2}=-b$.

### v2

@@ 3.1 Generalized Multivariate Bayesian Density-Density Regression
where $f$ is a measurable function that maps from $\mathbb{R}^{d_{1}}$ to $\mathbb{R}^{d_{2}}$, $w>0$, and $f\sharp F_{i}$ denotes the push-forward measure of $F_{i}$ through $f$.
@@
In (6) we are using SW distance to define a loss function that serves as generalized (negative log) likelihood. SW is an attractive choice for the loss function as it naturally defines a distance between random distributions $f\sharp{F}_{i}$ and ${G}_{i}$. One of the advantages of using the generalized likelihood is that it sidesteps the need to start with a full probabilistic description of all relevant unknowns. Let $Y_{i}=(y_{i1},\ldots,y_{im})$, when $G_{i}=\frac{1}{m}\sum_{j=1}^{m}\delta_{y_{ij}}$ for $i=1,\ldots,N$ is an empirical distribution with $m$ atoms, (6) implies a hypothetical sampling model
$Y_{i}\mid{F}_{i},f\sim p(Y_{i}\mid{F}_{i},f)\propto\exp\left(-wSW_{2}^{2}\left(f\sharp{F}_{i},\frac{1}{m}\sum_{j=1}^{m}\delta_{y_{ij}}\right)\right).$ (7)
Since $SW_{2}^{2}\left(f_{\sharp}F_{i},\frac{1}{m}\sum_{j=1}^{m}\delta_{y_{ij}}\right)$ is bounded due to the assumption of finite second moments of distributions, the normalizing constant in (7) is bounded. However, as a likelihood function of $f$ such normalization constant cannot be ignored, making (7) distinct from the generalized likelihood $\ell(f;G_{i},F_{i})$ in (6).
@@
To assess goodness of fit, we use the following residual error between the fitted distribution and the recorded response distributions: $PE(\mathcal{S})=\frac{1}{N}\sum_{i=1}^{N}SW_{2}^{2}(f\sharp{F}_{i},{G}_{i})$. Averaging with respect to a posterior sample $f_{1},\ldots,f_{T}\sim p(f\mid\mathcal{S})$, we have the following approximated mean predictive error: $\overline{\text{PE}}(\mathcal{S})=\frac{1}{T}\frac{1}{N}\sum_{t=1}^{T}\sum_{i=1}^{N}SW_{2}^{2}(f_{t}\sharp{F}_{i},{G}_{i}).$ To calibrate $\overline{\text{PE}}$ we introduce a reference model $f_{0}$ that serves as a worst-case reference (see later for examples), in a similar spirit as using an intercept-only model in a linear regression. We normalize the relative error to obtain the following relative residual error (RPE):
$\text{RPE}(\mathcal{S})=\frac{1}{N}\sum_{i=1}^{N}\frac{SW_{2}^{2}(f\sharp{F}_{i},{G}_{i})}{SW_{2}^{2}(f_{0}\sharp{F}_{i},{G}_{i})}.$ (8)
@@
When $\text{RPE}(\mathcal{S})=0$, it means a perfect fit as SW is a distance between distributions. When $\text{RPE}(\mathcal{S})=1$, it means a bad fit. Similarly, the mean RPE can be estimated using posterior samples, i.e.,
$\overline{\text{RPE}}(\mathcal{S})=\frac{1}{T}\frac{1}{N}\sum_{t=1}^{T}\sum_{i=1}^{N}\frac{SW_{2}^{2}(f_{t}\sharp{F}_{i},{G}_{i})}{SW_{2}^{2}(f_{0t}\sharp{F}_{i},{G}_{i})},$ (9)
@@
3.2 Posterior Consistency
The generalized likelihood (6) implies a generalized posterior
$p(f\mid\mathcal{S})\propto p(f)\exp\left(-w\sum_{i=1}^{N}SW_{2}^{2}(f\sharp{F}_{i},{G}_{i})\right),$ (10)
where $p(f)$ is a prior distribution of $f$. We now discuss the posterior consistency when $f$ takes a parametric form, i.e., $f_{\phi}$ with parameters $\phi\in\Phi$ and $(F_{1},G_{1}),\ldots,(F_{N},G_{N})\overset{i.i.d}{\sim}P$ for an unknown truth $P$. Under a prior $p(\phi)$, we have the generalized posterior:
$p(\phi\mid\mathcal{S})\propto p(\phi)\exp\left(-w\sum_{i=1}^{N}SW_{2}^{2}(f_{\phi}\sharp{F}_{i},{G}_{i})\right).$ (11)
Defining empirical risk and population risk
$R_{N}(\phi)=\frac{1}{N}\sum_{i=1}^{n}SW_{2}^{2}(f_{\phi}\sharp F_{i},G_{i}),\quad R(\phi)=\mathbb{E}_{(F,G)\sim P}[SW_{2}^{2}(f_{\phi}\sharp F,G)],$ (12)
we make the following assumptions for posterior consistency.
Assumption 1 (Identifiability).
There exists $\phi_{0}\in\Phi$ such that $R(\phi)$ attains its unique minimum at $\phi_{0}$. Moreover, for every $\epsilon>0$, $\Delta(\epsilon)\;=\;\inf_{\{\phi\in\Phi:\|\phi-\phi_{0}\|_{2}\geq\epsilon\}}\big(R(\phi)-R(\phi_{0})\big)\;>\;0.$
Assumption 2 (Compactness).
The parameter space $\Phi$ is compact.
Assumption 3 (Bounded second moments).
For all $(F,G)\in\mathrm{supp}(P)$, $\mathbb{E}_{X\sim F}[\|X\|_{2}^{2}]\leq M_{1},\quad\mathbb{E}_{Y\sim G}[\|Y\|_{2}^{2}]\leq M_{2},$ for some constants $M_{1},M_{2}<\infty$.
Assumption 4 (Prior positivity).
The prior $\pi$ assigns positive mass to every neighborhood of $\phi_{0}$; that is, for every $\epsilon>0$, $\pi\big\{B_{\epsilon}(\phi_{0})\big\}>0,\quad\text{where}\quad B_{\epsilon}(\phi_{0})=\{\phi\in\Phi:\|\phi-\phi_{0}\|_{2}<\epsilon\}.$
Assumption 5 (Regularity of Regression Function).
The regression function $f_{\phi}$ admits $\omega:\mathbb{R}_{+}\to\mathbb{R}_{+}$ ($\lim_{t\to 0}\omega(t)=0$) as a modulus of continuity: $\mathbb{E}_{X\sim F}\|f_{\phi}(X)-f_{\phi^{\prime}}(X)\|_{2}^{2}\leq\omega(\|\phi-\phi^{\prime}\|_{2}^{2})$ for all $(F,G)\in\mathrm{supp}(P)$ and $\phi,\phi^{\prime}\in\Phi$. For all $(F,G)\in\mathrm{supp}(P)$ and $\phi\in\Phi$, $\mathbb{E}_{X\sim F}[\|f_{\phi}(X)\|_{2}^{2}]\leq C,$ for a constant $C<\infty$.
Theorem 1.
Under Assumptions 15, for every $\epsilon>0$, the posterior measure $\pi_{N}$ of (11) satisfies
$\pi_{N}\big(\{\phi\in\Phi:\|\phi-\phi_{0}\|_{2}\geq\epsilon\}\big)\xrightarrow{a.s}0,$ (13)
as $N\to\infty$ under i.i.d sampling.
The proof of Theorem 1 is given in Appendix A.1, which requires proving the uniform law of a large number for $R_{N}(\phi)$ and $R(\phi)$.
3.3 Posterior Simulation
@@
We implement the posterior inference by Markov chain Monte Carlo (MCMC) simulation from (10). Again, we restrict the function $f$ to a parametric form, i.e., $f_{\phi}$ with parameters $\phi\in\Phi$ for computational efficiency. We restrict the function $f_{\phi}$ and the prior $p(\phi)$ to be differentiable with respect to $\phi$ to facilitate the use of gradient-based MALA (one-step hybrid Monte Carlo) transition probabilities using the following proposal distribution: $q(\phi^{\prime}\mid\phi)\propto\exp\left(-\frac{1}{4\eta}\|\phi^{\prime}-\phi-\eta\nabla_{\phi}\log p(\phi\mid\mathcal{S})\|_{2}^{2}\right)$, for a fixed step size $\eta>0$. We sample $\phi^{\prime}\sim q(\phi^{\prime}\mid\phi)$ as $\phi^{\prime}=\phi+\eta\nabla_{\phi}\log p(\phi\mid\mathcal{S})+\sqrt{2\eta}\epsilon_{0}$, with $\epsilon_{0}\sim\mathcal{N}(0,I)$, a standard multivariate Gaussian distribution of dimension matching the parameter $\phi$. We accept $\phi^{\prime}$ with probability: $\alpha=1\wedge{p(\phi\mid\mathcal{S})q(\phi\mid\phi^{\prime})}\big/{\left(p(\phi^{\prime}\mid\mathcal{S})q(\phi^{\prime}\mid\phi)\right)}.$
In practice, we cannot evaluate the generalized likelihood since $SW_{2}^{2}(f_{\phi}\sharp{F}_{i},{G}_{i})$ is intractable due to two reasons. First, we only observe $F_{i}$ and $G_{i}$ through their samples i.e., $x_{i1},\ldots,x_{im_{i1}}\sim F_{i}$ and $y_{i1},\ldots,y_{im_{i2}}\sim G_{i}$. Second, the expectation in the definition of SW (3) is intractable. As discussed in Section 2, we can use plug-in estimate (4) as a simple solution for the first issue. For the second issue, we can use Monte Carlo estimate as a solution (5). For the posterior inference purpose, we replace the generalized likelihood $p(\phi\mid\mathcal{S})$ by its estimation $\hat{p}(\phi\mid\mathcal{S})\propto p(\phi)\exp\left(-w\sum_{i=1}^{N}\widehat{SW}_{2}^{2}(f_{\phi}\sharp\widehat{F}_{i},\widehat{G}_{i};L)\right)$.
@@
Under these approximations we derive an estimate of the gradient of the log posterior density $\nabla_{\phi}\log p(\phi\mid\mathcal{S})\approx\nabla_{\phi}\log p(\phi)+\nabla\hat{p}(\mathcal{S}\mid\phi)=\nabla_{\phi}\log p(\phi)-w\sum_{i=1}^{N}\nabla_{\phi}\widehat{SW}_{p}^{p}(f_{\phi}\sharp\widehat{F}_{i},\widehat{G}_{i};L)$, where $\nabla_{\phi}\widehat{SW}_{2}^{2}(f_{\phi}\sharp\widehat{F}_{i},\widehat{G}_{i};L)=\frac{1}{L}\sum_{l=1}^{L}\nabla_{\phi}W_{2}^{2}(\theta_{l}\sharp f_{\phi}\sharp\widehat{F}_{i},\theta_{l}\sharp\widehat{G}_{i})$ with
@@
$\nabla_{\phi}W_{2}^{2}(\theta_{l}\sharp f_{\phi}\sharp\widehat{F}_{i},\theta_{l}\sharp\widehat{G}_{i})=\nabla_{\phi}\left(\min_{\gamma\in\Gamma(\mathbf{1}/m_{i1},\mathbf{1}/m_{i2})}\sum_{i^{\prime}=1}^{m_{i1}}\sum_{j^{\prime}=1}^{m_{i2}}(\theta_{l}^{\top}f_{\phi}(x_{ii^{\prime}})-\theta_{l}^{\top}y_{ij^{\prime}})^{2}\gamma_{i^{\prime}j^{\prime}}\right).$ (14)
@@
$\nabla_{\phi}\left(\min_{\gamma\in\Gamma\left(\frac{\mathbf{1}}{m_{i1}},\frac{\mathbf{1}}{m_{i2}}\right)}\sum_{i^{\prime}=1}^{m_{i1}}\sum_{j^{\prime}=1}^{m_{i2}}(\theta^{\top}f_{\phi}(x_{ii^{\prime}})-\theta^{\top}y_{ij^{\prime}})^{2}\gamma_{i^{\prime}j^{\prime}}\right)\\ =\nabla_{\phi}\left(\sum_{i^{\prime}=1}^{m_{i1}}\sum_{j^{\prime}=1}^{m_{i2}}(\theta^{\top}f_{\phi}(x_{ii^{\prime}})-\theta^{\top}y_{ij^{\prime}})^{2}\gamma_{i^{\prime}j^{\prime}}^{\star}\right)=\sum_{i^{\prime}=1}^{m_{i1}}\sum_{j^{\prime}=1}^{m_{i2}}\gamma_{i^{\prime}j^{\prime}}^{\star}\nabla_{\phi}(\theta^{\top}f_{\phi}(x_{ii^{\prime}})-\theta^{\top}y_{ij^{\prime}})^{2}\\ =2\sum_{i^{\prime}=1}^{m_{i1}}\sum_{j^{\prime}=1}^{m_{i2}}\gamma_{i^{\prime}j^{\prime}}^{\star}(\theta^{\top}f_{\phi}(x_{ii^{\prime}})-\theta^{\top}y_{ij^{\prime}})\nabla_{\phi}(\theta^{\top}f_{\phi}(x_{ii^{\prime}})).$ (15)
The calculation of $\nabla_{\phi}(\theta^{\top}f_{\phi}(x_{ii^{\prime}}))$ depends on the form of $f_{\phi}$.
In the upcoming examples and simulations, we utilize a linear function for simplicity, $f_{A,b}(x)=Ax+b$, i.e., $\phi=(A,b)$ with $A=[A_{ij}]\in\mathbb{R}^{d_{2}\times d_{1}}$ and $b\in\mathbb{R}^{d_{2}}$. It is easy to check that this function satisfies Assumption 5, i.e., it is Lipchitz in $\phi$ (see Proposition 1 in Appendix A.1). While any alternative functions can be used e.g., deep neural network (Wilson and Izmailov,, 2020), we choose a linear form as the most parsimonious model, and recommend more complex choices only if residual errors so indicate. We put a horseshoe prior (Carvalho et al.,, 2009) on $A$, which will be later discussed in detail in the context of graph discovery:
@@
and $b\sim\mathcal{N}(0,I)$. We implement posterior simulation using standard MCMC transition probabilities using the inverse gamma complete conditional posterior distributions for $\lambda_{ij}^{2}$, $\nu_{ij}$, $\tau^{2}$, and $\zeta$. See Supplementary Materials  B for more details of the MCMC sampler. To sample from $p(A,b\mid\lambda^{2},\tau^{2},\mathcal{S})$, we use the MALA algorithm as discussed above. Substituting in $\nabla_{A}(\theta^{\top}(Ax_{ii^{\prime}}+b))=\theta x_{ii^{\prime}}^{\top}$ and $\nabla_{b}(\theta^{\top}(Ax_{ii^{\prime}}+b))=\theta$ into Equation (15), we obtain the gradient for the log-likelihood. For the gradient of the log-prior, we have $\nabla_{A_{ij}}\log p(A_{ij}\mid\lambda_{ij}^{2},\tau^{2})=\nabla_{A_{ij}}\frac{-A_{ij}^{2}}{2\lambda_{ij}^{2}\tau^{2}}=\frac{-A_{ij}}{\lambda_{ij}^{2}\tau^{2}}$ and $\nabla_{b}\log p(b)=\nabla_{b}\frac{-\|b\|_{2}^{2}}{2}=-b$.